OpenStudy (anonymous):

Find the next three terms of the arithmetic sequence. Then, give the common difference, and write a rule to describe the sequence. 1. 66, 77, 88, 99, . . . (0 points) Tell whether each sequence is arithmetic , geometric , or neither . Then find the next three terms. For each arithmetic or geometric sequence, give the common difference or ratio, and write a rule to describe the sequence. 2. 9, 3, 1, onethird_large, . . . (0 points) 3. 1, 4, 9, 16, . . . (0 points)

OpenStudy (anonymous):

1. Arithmetic sequence meaning, thats right it means sequences of numbers that follow a pattern of adding a fixed number from one term to the next 66 to 77 add 11, 77 to 88 add 11, 99 to 110 add 11. n = next term, p = previous term. n = p + 11 2. Sequences of numbers that follow a pattern of multiplying a fixed number from one term to the next are called geometric sequences. Therefore 9* 1/3 = 3, 3 * 1/3 = 1 and so on. 3. Neither